OpenStudy (anonymous):

Where 0<=x<360, Solve the equation. 3sec(x)-8=0

geerky42 (geerky42):

\[3\sec x = 8\]\[\sec x = \dfrac{8}{3}\]\[\dfrac{1}{\cos x} = \dfrac{8}{3}\]\[\cos x = \dfrac{3}{8}\]

OpenStudy (anonymous):

If x={68,292}, is there any way to do it without a calculator?

OpenStudy (tkhunny):

Well, you CAN construct a 17-gon with a straight edge and compass, but why? Use the calculator. Note: Are you SURE you want to go with no decimal places?